In a typical parallel plate capacitor the separati

I mostly just need to confirm my answers thanks!
Answer #1

2.1 For a parallel plate capacitor, C = Ae/d and Q = CV
so, when d is increased, C decreases
and Q is constasnt
so, V increases
so, b) V increasesm but E = V/d = constant
2.2 wo wires carrying current in same direction always attract each other. SO the current is a) parallel
2.3 The proton will have the force acting on the opposite direction, due to opposite charge
also, F = qVB
and F = mv^2/r
so mv/r = qB
r = mv/qB
so mass of proton is more, so more radius
hence
d) the proton turns in opposite irection with larger radius
2.4 The component of the velocity parallel to field remains constant, the component perpendicular to the fiedld causes a force ont he particvel causing it to move in circular direction
hence the net motion is a helix